# A modified recursive transfer matrix algorithm for radiation and scattering computation of multilayer spheres

### v1

@@ 4 Results
The following section presents three test cases using the algorithm proposed in this paper. The first test case examines the scattering problem of a double-layered sphere with varying radii, where the refractive indices of the inner and outer layers are specified and compared with the RTMA. The second test case involves scattering by a multilayered sphere with the same radius, where the refractive indices of each shell are randomly generated, and the extinction cross-section is calculated for spheres with various numbers of layers. The third test case addresses the thermal radiation problem of a multilayered sphere, calculating the emissivity of a double-layered sphere.
@@
In the first test, a coated sphere is considered, with the refractive indices of the core and shell layers set to $m_{1}=1.33+i0.0$ and $m_{2}=1.33+i$, respectively, and the core radius accounting for 50% of the total radius. The extinction efficiency $Q_{ext}$ is calculated for size parameters $x\in[1.0,1200.0]$. The RTMA experiences numerical overflow for $x>700.0$, while the mRTMA remains stable, as shown in Figure 2.
@@
In the second test, we examined the scattering of a multilayered sphere with randomly generated refractive indices, with the total number of layers ranging from 2 to 2000. The refractive indices are generated as $n=n^{\prime}+in^{\prime\prime}$, with $n^{\prime}\sim\mathcal{U}_{[1,2]}$ and $\log_{10}n^{\prime\prime}\sim\mathcal{U}_{[-3,1]}$, where $\mathcal{U}$ denotes a uniform distribution. As shown in Figure 3, the mRTMA demonstrates significant improvement over the RTMA.
@@
In the third test, the mRTMA is used to calculate the electromagnetic radiation of a double-layered sphere with varying radii, with the sphere maintained at a temperature of 300K in a cold environment. Figure 4 shows the radiation characteristics of the double-layered sphere, using SiC and gold as example materials, representing a dielectric and a conductor, respectively.
@@ 5 Conclusion
This paper reviews the Lorentz-Mie theory and its development in the context of electromagnetic scattering by multilayered spheres. It explores the numerical stability issues associated with multilayered sphere scattering and traces the evolution of various improved algorithms. By summarizing and comparing different recursive methods and transfer matrix algorithms, we identify the limitations and potential issues of these approaches, particularly when dealing with multilayered spheres, thin shells, and strongly absorbing media. Building on these earlier works, we propose an efficient recursive transfer matrix algorithm, demonstrating that it achieves high accuracy over a broad range of size parameters and significantly extends the applicability compared to the classical transfer matrix methods. Additionally, we propose and extend an algorithm for calculating spherical Bessel functions. Future work can focus on optimizing these numerical methods to further enhance their stability and applicability, providing more effective tools for solving complex electromagnetic scattering problems.

### v2

@@ 4 Results
The following section presents three test cases using the algorithm proposed in this paper. The first test case examines the scattering problem of a double-layered sphere with varying radii, where the refractive indices of the inner and outer layers are specified and compared with the RTMA. The second test case involves scattering by a multilayered sphere with the same radius, where the refractive indices of each shell are randomly generated, and the extinction cross-section is calculated for spheres with various numbers of layers. The third test case addresses the thermal radiation problem of a multilayered sphere, calculating the emissivity of a double-layer sphere.
Figure 2: (a) Plot of the extinction efficiency $Q_{ext}$ for the coated sphere as a function of the outermost layer’s size parameter. (b) Loglog plot of $Q_{ext}$ - 2 of the coated sphere as a function of the outermost layer’s size parameter. The plot compares the results with those obtained using RTMA. The refractive indices of the core and shell are $m_{1}=1.33+0.0\mathrm{i}$ and $m_{2}=1.33+\mathrm{i}$, respectively.
@@
In the 1st test, a coated sphere is considered, with the refractive indices of the core and shell layers set to $m_{1}=1.33+0.0\mathrm{i}$ and $m_{2}=1.33+\mathrm{i}$, respectively, and the core radius accounting for 50% of the total radius. The extinction efficiency $Q_{ext}$ is calculated for size parameters $x\in[1.0,1200.0]$. The RTMA experiences numerical overflow for $x>700.0$, while the mRTMA remains stable, as shown in Figure 2.
Figure 3: (a) The extinction efficiency $Q_{\rm ext}$ of the multilayer sphere as a function of the number of layers. The refractive index of each layer is randomly generated. (b) Absolute difference between $Q_{\rm ext}$ and $Q_{\rm ext}^{\rm Mie}$, where $Q_{\rm ext}^{\rm Mie}$ is the extinction efficiency for a homogeneous sphere, computed using the Mie algorithm [20]. The refractive index is fixed at $1.33+\mathrm{i}$. The size parameter of the sphere in (a) and (b) is $4\pi$.
@@
In the 2nd test, we examined the scattering of a multilayered sphere with randomly generated refractive indices, with the total number of layers ranging from 2 to 2002. The refractive indices are generated as $n=n^{\prime}+\mathrm{i}n^{\prime\prime}$, with $n^{\prime}\sim\mathcal{U}_{[1,2]}$ and $\log_{10}n^{\prime\prime}\sim\mathcal{U}_{[-3,1]}$, where $\mathcal{U}$ denotes a uniform distribution. Additionally, we also tested the case of a multilayered sphere where all layers have the same refractive index of $1.33+\mathrm{i}$, and compared the results with the Mie scattering results for a homogeneous sphere. As shown in Figure LABEL:fig:multi, the mRTMA demonstrates significant improvement over the RTMA.
Figure 4: The emitted energy of a SiC+gold coated sphere at a temperature of 300K, as a function of the outer radius R, is normalized with respect to the Stefan-Boltzmann result. Different curves correspond to different ratios of inner to outer layer radii. For SiC and gold, the optical constants’ data are from [21] and [22] respectively.
@@
In the 3rd test, the mRTMA is used to calculate the electromagnetic radiation of a double-layer sphere with varying radii, with the sphere maintained at a temperature of 300K in a cold environment. Figure 4 shows the radiation characteristics of the double-layer sphere, using SiC and gold as example materials, representing a dielectric and a conductor, respectively.
Figure 5: Performance validation. Prog. 1: Our implementation of the mRTMA. Prog. 2: Our implementation of Yang’s recursive algorithm. Prog. 3: Scattnlay’s implementation of Yang’s recursive algorithm. Size parameter of the sphere is $4\pi$.
In the 4th test, we validate the precision of mRTMA by comparing it with RTMA and Yang’s recursive algorithm. The multilayered sphere is simulated using randomly generated refractive indices, with the total number of layers ranging from 2 to 702, increasing in steps of 50 layers. The refractive indices are generated as $n=n^{\prime}+\mathrm{i}n^{\prime\prime}$, with $n^{\prime}\sim\mathcal{U}_{[1,2]}$ and $\log_{10}n^{\prime\prime}\sim\mathcal{U}_{[-3,1]}$, where $\mathcal{U}$ denotes a uniform distribution. As shown in Figure 5, the RTMA, mRTMA, and Yang’s recursive algorithm exhibit excellent agreement with high precision in this case, demonstrating the validity of the mRTMA.
@@ 5 Conclusion
This paper aimed to develop a new, efficient algorithm for electromagnetic radiation and scattering by a multilayered sphere based on Lorentz-Mie theory, with a particular emphasis on addressing numerical stability issues. We examined the evolution of various improved algorithms, summarizing and comparing recursive methods and transfer matrix algorithms to identify their limitations and potential challenges. These challenges are especially pronounced in scenarios involving thin shells and strongly absorbing layers.
Building on earlier works, we proposed an efficient recursive transfer matrix algorithm (mRTMA). The primary contribution of mRTMA lies in the development of new recursive formulas based on the Transfer Matrix Method (RTMA). These formulas effectively address the numerical overflow issues encountered in RTMA and enable the algorithm to handle multilayer spheres embedded in absorptive host media. Like RTMA, mRTMA can simultaneously compute the expansion coefficients for spherical layers.
Moreover, by adopting a hybrid recursive strategy, mRTMA resolved the singularity problem associated with logarithmic derivatives in Yang’s algorithm. In addition, unlike both RTMA and Yang’s algorithm, mRTMA introduced recursive formulas specifically designed for the outgoing wave scenario, where the incident wave originates from within the sphere. These formulas broaden the applicability of the method to specific scenarios, such as spherical cavities and spherical antennas.
Finally, we demonstrated that mRTMA achieves high accuracy across a broad range of size parameters and significantly extends the applicability of classical transfer matrix methods. Future work can focus on optimizing these numerical methods to further enhance their stability and efficiency, providing more effective tools for solving complex electromagnetic scattering problems.